OpenStudy (anonymous):

What would this be classified as and how do we solve this? 2(m − 4) + 1?

OpenStudy (anonymous):

@undeadknight26

OpenStudy (anonymous):

It is classified as Expression...

OpenStudy (anonymous):

How do we solve this?

OpenStudy (anonymous):

I know Pemdas

OpenStudy (anonymous):

But I need step by step on this example

undeadknight26 (undeadknight26):

I come and all the good guys show up the solve the equation lel

OpenStudy (anonymous):

Woah! its the nerd squad !

undeadknight26 (undeadknight26):

LOL the hot nerd squad B-) The nerd squad with fast and loud cars B-)

OpenStudy (anonymous):

Yeah... Hotwheels 4tw

OpenStudy (anonymous):

I knew my car collection was gonna come in handy some day!

OpenStudy (anonymous):

Water you still typin'

OpenStudy (anonymous):

Yeah I am still typing...

OpenStudy (anonymous):

Use Distributive Property: \[a(b+c) = ab + ac \\ m(n-q) = mn - {mq}\]

OpenStudy (anonymous):

Firstly you should distribute \(2\) into the brackets like I did above..

OpenStudy (anonymous):

How

OpenStudy (anonymous):

@DarkMoonZ